### v1

# A Heuristic ADMM-based Approach for Tree-Constrained Optimization

Abstract
This paper presents centralized and distributed Alternating Direction Method of Multipliers (ADMM) frameworks for solving large-scale nonconvex optimization problems with binary decision variables subject to spanning tree or rooted arborescence constraints. We address the combinatorial complexity by introducing a continuous relaxation of the binary variables and enforcing agreement through an augmented Lagrangian formulation. The algorithms alternate between solving a convex continuous subproblem and projecting onto the tree-feasible set, reducing to a Minimum Spanning Tree or Minimum Weight Rooted Arborescence problem, both solvable in polynomial time. The distributed algorithm enables agents to cooperate via local communication, enhancing scalability and robustness. We apply the framework to multicommodity flow design with hop-constrained spanning trees. Numerical experiments demonstrate that our methods yield high-quality feasible solutions, achieving near-optimal performance with significant computational savings compared to the commercial solver Gurobi.
Index Terms:
Networks, distributed algorithms, consensus, optimization
I Introduction
I-A The Problem
Let $(\mathcal{V},\mathcal{E})$ (resp. $(\mathcal{V},\mathcal{A})$) be an undirected (respectively directed) graph, where $\mathcal{V}$ denotes the set of vertices and $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ (respectively $\mathcal{A}\subset\mathcal{V}\times\mathcal{V}$) denotes the set of edges (resp. arcs). Throughout this work, we denote $|\mathcal{V}|=n$ and $|\mathcal{E}|=m$ (resp. ($|\mathcal{A}|=m$)). Bold symbols consistently represent vectors or matrices; for example, if $\boldsymbol{x}\in\mathbb{R}^{m}$, then $\boldsymbol{x}=(x_{1},x_{2},...,x_{m})^{T}$.
We consider the following Mixed-Integer Non-linear Programming (MINLP) problem:
$\begin{array}[]{ll}\text{minimize}&f\left(\boldsymbol{x},\boldsymbol{z}\right)\\[3.00003pt] \text{subject to}&g\left(\boldsymbol{x},\boldsymbol{z}\right)\leq\boldsymbol{0}_{q},\\[1.99997pt] &\left(\boldsymbol{x},\boldsymbol{z}\right)\in\mathcal{X}\times\mathcal{Z},\end{array}$ (1)
where the continuous decision variable $\boldsymbol{x}\in\mathbb{R}^{m}$ belongs to the set $\mathcal{X}\subseteq\mathbb{R}^{m}$, and $\boldsymbol{z}$ is the binary decision variable in $\mathcal{Z}$ indicating the activation status of the edges or arcs; that is, $z_{ij}=1$ if the edge or arc $(i,j)\in\mathcal{E}$ or $\mathcal{A}$ is active, and $z_{ij}=0$ otherwise. The number of constraints is denoted by $q$.
A topological constraint is imposed on $\boldsymbol{z}$, requiring it to define either a spanning tree (in the undirected case) or a rooted arborescence (in the directed case) with a designated root node. Formally,
$\mathcal{Z}=\left\{\begin{array}[]{c}\boldsymbol{z}\in\{0,1\}^{m}:\boldsymbol{z}\ \text{induces a spanning tree}\\ \text{ or a rooted arborescence}\end{array}\right\}.$
We assume that both the objective function $f:\mathcal{X}\times\tilde{\mathcal{Z}}\to\mathbb{R}$ and the constraint function $g:\mathcal{X}\times\tilde{\mathcal{Z}}\to\mathbb{R}^{q}$ are jointly convex in their continuous arguments:
$(\boldsymbol{v},\boldsymbol{w})\mapsto f(\boldsymbol{v},\boldsymbol{w})\quad\text{and}\quad(\boldsymbol{v},\boldsymbol{w})\mapsto g(\boldsymbol{v},\boldsymbol{w}),$
are convex over $\mathcal{X}\times\tilde{\mathcal{Z}}$, where $\tilde{\mathcal{Z}}=[0,1]^{m}$ denotes the convex hull of $\mathcal{Z}$.
I-B Motivation and Applications
Problem (1) addresses the design of network topologies that must satisfy both continuous operational constraints and discrete structural requirements; specifically, that the selected edges in the undirected case form a spanning tree, or, in the directed case, a rooted arborescence. In many real-world applications, additional structural and operational constraints further increase the problem’s complexity, leading to exponential growth with respect to the network size.
For example, enforcing a maximum diameter or hop limit on the tree possibly in conjunction with multi-commodity flow requirements is essential in telecommunication and transportation networks, where bounded path lengths ensure acceptable latency and service quality [1, 2, 3, 4]. In electric power systems, radial spanning-tree constraints are fundamental in distribution network reconfiguration problems, where the objective is to minimize power losses, balance loads, or improve reliability while preserving a radial topology for protection coordination and safe operation [5, 6, 7, 8, 9, 10]. These diverse applications underscore the need for scalable algorithms capable of handling both centralized and distributed settings on directed or undirected graphs solving problem (1).
I-C Solution Methods
Problem (1) is a MINLP, for which both exact and approximate solution strategies exist. Among the exact approaches, the most direct method is complete enumeration, in which all feasible combinations of the discrete variables are systematically examined, the corresponding convex optimization subproblem is solved to optimality, and the configuration with the smallest objective value is retained.
More sophisticated global methods include branch-and-bound [11, 12] and branch-and-cut [13, 14], which guarantee identification of the global optimum. Cutting-plane techniques [15, 16] iteratively solve relaxations of the original problem and add linear inequalities to progressively enforce integrality. While these methods provide theoretical guarantees, they typically have non-polynomial worst-case complexity, making them impractical for large-scale or embedded optimization, and their runtime can vary significantly across instances.
Heuristic methods can quickly produce high-quality, though not necessarily optimal, solutions. Examples include the relax-and-round approach, which solves a convex relaxation and then projects the result onto the original nonconvex set, and approaches that fix nonconvex variables to plausible values before solving the remaining convex subproblems [17]. Feasibility-oriented heuristics, such as the feasibility pump [18, 19], aim to find feasible solutions efficiently. Although lacking theoretical guarantees, these methods are often effective in practice and well-suited to time-constrained settings.
I-D The Alternating Direction Method
The Alternating Direction Method of Multipliers (ADMM) is a primal–dual splitting algorithm originally developed for convex optimization problems [20, 21]. Its use as a general heuristic for nonconvex optimization has been explored in, e.g., [20, Ch. 9] and [22], and it has recently attracted attention as a practical approach for obtaining approximate solutions to NP-hard problems. ADMM can address computationally challenging MINLPs [17] when projection onto the discrete constraint set can be performed, either exactly or approximately. While global convergence is not guaranteed in the nonconvex setting, ADMM often yields high-quality local solutions.
ADMM has been successfully applied in various mixed-integer optimization contexts, including mixed-integer quadratic programming [23], pump scheduling and water network management [24], weighted network design with cardinality constraints [25], and electric distribution system reconfiguration [6, 9, 10]. Its main strength lies in decomposing large-scale problems into simpler subproblems that can be solved sequentially while maintaining global coordination. This makes ADMM a scalable tool for centralized optimization tasks and a natural choice for designing distributed optimization algorithms that converge under mild assumptions [20, 26, 27, 28, 29].
I-E Distributed Optimization
Distributed optimization encompasses a class of algorithms in which multiple agents cooperate to solve a global optimization problem. Such methods are valued for their scalability, robustness, privacy preservation, and adaptability. By partitioning the computational workload among agents, these algorithms are well-suited to large-scale systems, can tolerate local faults or communication losses, and alleviate centralized bottlenecks. For privacy-sensitive scenarios, distributed schemes enable local data processing without sharing raw information, thereby reducing both privacy risks and communication overhead compared to fully centralized architectures [27, 26].
Applications span a wide range of domains, including networked multi-agent coordination, large-scale machine learning (e.g., distributed training of Support Vector Machines), and smart grid management [30, 31, 27, 32, 10]. A canonical formulation involves $N$ agents jointly solving
$\underset{x\in\mathbb{R}^{N}}{\min}\ \sum_{i=1}^{N}f_{i}(x),$ (2)
where $f_{i}:\mathbb{R}^{N}\to\mathbb{R}$ denotes the local objective of agent $i$ and $x$ is a global decision variable shared by all agents.
Although distributed optimization is well-developed for convex and continuous problems [33], many practical settings, such as network reconfiguration, facility location, and scheduling, necessitate mixed-integer formulations with binary or general discrete variables. These distributed MINLPs are substantially more difficult due to their combinatorial nature, non-convex feasible regions, and the absence of scalable, exact distributed solvers [34, 35]. In such contexts, enforcing global combinatorial constraints (e.g., radiality in network topologies) while relying only on local agent knowledge typically demands consensus-based or primal–dual decomposition methods, augmented with advanced projection or relaxation strategies.
I-F Related Works and Contribution
In the context of heuristic methods for optimization problems with tree constraints, several works have proposed distributed ADMM-based approaches to address the power flow network reconfiguration problem under radiality constraints [6, 7, 8]. However, these methods do not guarantee that the final solution satisfies the tree constraint. The main limitation lies in the projection step, which often relies on rounding of relaxed continuous variables to binary decisions, without explicitly enforcing the radiality condition. As a result, the obtained topology may contain cycles or be disconnected, violating feasibility.
This limitation has been addressed in subsequent works in both centralized [9] and distributed [10] settings, where the projection step is reformulated as an exact combinatorial optimization problem. Specifically, as an MST problem in the undirected case or a MWRA problem in the directed case [36, Chapter 6.3]. This guarantees that the resulting topology is always a feasible tree structure, thereby ensuring strict satisfaction of the radiality constraint.
In this work, we extend prior ADMM-based approaches [6, 7, 8] by developing centralized and distributed algorithms that ensure tree constraints via exact MST or MWRA projections, applicable to directed and undirected graphs. As a case study, we demonstrate the effectiveness of both algorithms on multicommodity flow formulations for spanning trees subject to hop constraints [1, 2, 3, 4].
II The Algorithm
II-A Review
Before presenting the proposed algorithms, we recall two fundamental combinatorial optimization problems:
II-A1 Minimum Spanning Tree (MST)
The MST problem arises in undirected graphs and provides the basis for enforcing tree structures in $\mathcal{Z}$ for undirected network design. Let $G=(\mathcal{V},\mathcal{E})$ be an undirected graph, and let $\{h_{ij}\}_{(i,j)\in\mathcal{E}}$ denote the set of edge weights. The MST problem seeks a subset of edges $\mathcal{T}\subset\mathcal{E}$ such that $(\mathcal{V},\mathcal{T})$ forms a spanning tree, i.e., a connected and acyclic subgraph that includes all vertices, while minimizing the total weight:
$\sum_{(i,j)\in\mathcal{T}}h_{ij}.$ (3)
The MST problem can be solved in polynomial time via classical algorithms such as Kruskal’s or Prim’s method, with respective time complexities $O(m\log n)$ and $O(m+n\log n)$ [36, Chapter 6.3].
II-A2 Minimum Weight Rooted Arborescence (MWRA)
The MWRA problem generalizes the MST concept to directed graphs. Given a directed graph $G=(\mathcal{V},\mathcal{A})$, a designated root node $r\in\mathcal{V}$, and a set of arc weights $\{h_{ij}\}_{(i,j)\in\mathcal{A}}$, the goal is to identify a directed spanning tree (arborescence) $\mathcal{T}\subset\mathcal{A}$ rooted at $r$ such that there exists a unique directed path from $r$ to every other node, while minimizing the total weight:
$\sum_{(i,j)\in\mathcal{T}}h_{ij}.$ (4)
This problem can be solved in polynomial time using Edmonds’ algorithm, with complexity $O(mn)$ [36, Chapter 6.3].
II-B The centralized ADMM Algorithm
To address the combinatorial nature of the problem, we introduce a continuous relaxation variable $\boldsymbol{w}\in[0,1]^{m}$ to serve as a surrogate for the binary variable $\boldsymbol{z}$. This yields the following relaxed formulation:
$\begin{array}[]{ll}\text{minimize}&f(\boldsymbol{x},\boldsymbol{w})\\ \text{subject to}&g(\boldsymbol{x},\boldsymbol{w})\leq\boldsymbol{0}_{q},\\ &\boldsymbol{w}=\boldsymbol{z},\\ &(\boldsymbol{x},\boldsymbol{w},\boldsymbol{z})\in\mathcal{X}\times[0,1]^{m}\times\mathcal{Z}.\end{array}$
The corresponding augmented Lagrangian function is defined as:
$\mathcal{L}_{\rho}(\boldsymbol{x},\boldsymbol{w},\boldsymbol{z},\boldsymbol{\mu})=f(\boldsymbol{x},\boldsymbol{w})+\boldsymbol{\mu}^{\top}(\boldsymbol{z}-\boldsymbol{w})+\frac{\rho}{2}\left\|\boldsymbol{z}-\boldsymbol{w}\right\|^{2},$
where $\boldsymbol{\mu}\in\mathbb{R}^{m}$ is the vector of Lagrange multipliers, and $\rho>0$ is a penalty parameter.
We define the set
$\Sigma=\left\{(\boldsymbol{x},\boldsymbol{w})\in\mathcal{X}\times[0,1]^{m}:g(\boldsymbol{x},\boldsymbol{w})\leq\boldsymbol{0}_{q}\right\}.$
and we assume that it is nonempty. The ADMM update steps then read:
$(\boldsymbol{x},\boldsymbol{w})_{k+1}$ $=\underset{(\boldsymbol{x},\boldsymbol{w})\in\Sigma}{\arg\min}\left[f(\boldsymbol{x},\boldsymbol{w})+\frac{\rho}{2}\|\boldsymbol{z}_{k}-\boldsymbol{w}+\boldsymbol{\mu}_{k}\|^{2}\right],$ (5a)
$\boldsymbol{z}_{k+1}$ $=\underset{\boldsymbol{z}\in\mathcal{Z}}{\arg\min}\|\boldsymbol{z}-\boldsymbol{w}_{k+1}+\boldsymbol{\mu}_{k}\|^{2},$ (5b)
$\boldsymbol{\mu}_{k+1}$ $=\boldsymbol{\mu}_{k}+\boldsymbol{z}_{k+1}-\boldsymbol{w}_{k+1}.$ (5c)
The subproblem (5a) is a continuous convex optimization problem that can be efficiently solved using standard convex optimization solvers. In contrast, problem (5b) is a MINLP, specifically an integer quadratic programming problem, whose solution enforces a spanning tree or rooted arborescence structure. In this case, the step reduces to solving an MST problem or MWRA problem with iteratively updated edge weights.
Proposition II.1
Problem (5b) is equivalent to solving the following discrete optimization problem at each iteration:
$\boldsymbol{z}_{k+1}=\underset{\boldsymbol{z}\in\mathcal{Z}}{\arg\min}\;\boldsymbol{z}^{T}\boldsymbol{h}_{k},$ (6)
where the weight vector $\boldsymbol{h}_{k}\in\mathbb{R}^{m}$ is given by
$\boldsymbol{h}_{k}=\boldsymbol{\mu}_{k}-\boldsymbol{w}_{k+1}.$ (7)
Proof:
We begin by expanding the squared norm appearing in the second ADMM subproblem: We aim to show that problem (5b) is equivalent to solving the discrete optimization problem (6). Starting with the objective in (5b), we expand the squared norm:
$\left\|\boldsymbol{z}-\boldsymbol{w}_{k+1}+\boldsymbol{\mu}_{k}\right\|^{2}$ $=$ $\boldsymbol{z}^{\top}\boldsymbol{z}+\left\|\boldsymbol{\mu}_{k}-\boldsymbol{w}_{k+1}\right\|^{2}$
$+2\boldsymbol{z}^{\top}\left(\boldsymbol{\mu}_{k}-\boldsymbol{w}_{k+1}\right).$
Since $\boldsymbol{z}\in\mathcal{Z}\subset\{0,1\}^{m}$ represents a spanning tree or arborescence with exactly n1n-1 edges, we have $\boldsymbol{z}^{\top}\boldsymbol{z}=\sum_{(i,j)\in\mathcal{E}\text{ or }\mathcal{A}}z_{ij}=n-1$, as each selected edge contributes 1 to the sum. Thus, the expression becomes:
$\left\|\boldsymbol{z}-\boldsymbol{w}_{k+1}+\boldsymbol{\mu}_{k}\right\|^{2}$
$=$ $(n-1)+\left\|\boldsymbol{\mu}_{k}-\boldsymbol{w}_{k+1}\right\|^{2}+2\boldsymbol{z}^{\top}\left(\boldsymbol{\mu}_{k}-\boldsymbol{w}_{k+1}\right)$
$=$ $(n-1)+\left\|\boldsymbol{\mu}_{k}-\boldsymbol{w}_{k+1}\right\|^{2}+2\boldsymbol{z}^{\top}\boldsymbol{h}_{k}.$
Since n1n-1 and $\left\|\boldsymbol{\mu}_{k}-\boldsymbol{w}_{k+1}\right\|^{2}$ are constant with respect to $\boldsymbol{z}$, minimizing the expression over $\boldsymbol{z}\in\mathcal{Z}$ is equivalent to minimizing $\boldsymbol{z}^{\top}\boldsymbol{h}_{k}$. Thus, problem  (5b) reduces to:
$\boldsymbol{z}_{k+1}=\underset{\boldsymbol{z}\in\mathcal{Z}}{\arg\min}\ \boldsymbol{z}^{\top}\boldsymbol{h}_{k},$
which matches (6). ∎
Problem (6) is an integer linear optimization problem. When the set $\mathcal{Z}$ corresponds to the collection of indicator vectors of spanning trees in an undirected graph, it reduces to the MST problem (3). Conversely, if $\mathcal{Z}$ represents the set of rooted arborescences in a directed graph, it becomes the MWRA problem (4). In both formulations, the edge or arc weights are determined by the components of the vector $\boldsymbol{h}_{k}$ at each iteration $k\geq 1$.
II-C The Distributed ADMM Algorithm
We assume that the objective function $f$ and the constraint function $g$ are separable, i.e.,
$f(\boldsymbol{x},\boldsymbol{z})=\sum_{i\in\mathcal{V}}f^{i}(\boldsymbol{x},\boldsymbol{z}),\quad\sum_{i\in\mathcal{V}}g^{i}(\boldsymbol{x},\boldsymbol{z})\leq\boldsymbol{0}_{q},$
where $f^{i}$ and $g^{i}$ denote, respectively, the local objective function and the local constraint vector associated with agent $i\in\mathcal{V}$. The distributed formulation of problem (8) can thus be expressed as:
$\begin{array}[]{ll}\text{minimize}&\sum_{i\in\mathcal{V}}f^{i}(\boldsymbol{x},\boldsymbol{z})\\ \text{subject to}&\sum_{i\in\mathcal{V}}g^{i}(\boldsymbol{x},\boldsymbol{z})\leq\boldsymbol{0}_{q},\\ &(\boldsymbol{x},\boldsymbol{z})\in\mathcal{X}\times\mathcal{Z}.\end{array}$ (8)
Consider a directed graph $G=(\mathcal{V},\mathcal{A})$. Our objective is to establish a distributed method that enables the agents to cooperatively solve the optimization problem (8) while exchanging information only with their neighbors in $G$.
Fix any $\rho>0$ and initialize, for each agent $i\in\mathcal{V}$, the dual states as $\boldsymbol{\nu}^{\,i}_{0}=\boldsymbol{\xi}^{\,i}_{0}=\boldsymbol{0}_{m}$. Then, for every iteration $k\in\mathbb{N}$, each agent $i\in\mathcal{V}$ performs the following update steps:
$(\boldsymbol{x}^{i},\boldsymbol{w}^{i})_{k+1}=\underset{(\boldsymbol{x}^{i},\boldsymbol{w}^{i})\in\Sigma^{i}}{\arg\min}\left\{f^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i})\right$.
$+\frac{\rho}{2}\left\|\boldsymbol{z}_{k}^{i}-\boldsymbol{w}^{i}+\boldsymbol{\mu}_{k}^{i}\right\|^{2}+(\boldsymbol{\nu}_{k}^{i})^{\top}\boldsymbol{x}^{i}+(\boldsymbol{\xi}_{k}^{i})^{\top}\boldsymbol{w}^{i}$
$+\frac{\rho}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left\|\boldsymbol{x}^{i}-\frac{\boldsymbol{x}_{k}^{i}+\boldsymbol{x}_{k}^{j}}{2}\right\|^{2}$
$+\left.\frac{\rho}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left\|\boldsymbol{w}^{i}-\frac{\boldsymbol{w}_{k}^{i}+\boldsymbol{w}_{k}^{j}}{2}\right\|^{2}\right\},$
$\boldsymbol{z}_{k+1}^{i}$ $=$ $\underset{\boldsymbol{z}^{i}\in\mathcal{Z}}{\arg\min}\left\|\boldsymbol{z}^{i}-\boldsymbol{w}_{k+1}^{i}+\boldsymbol{\mu}_{k}^{i}\right\|^{2},$ (9b)
$\boldsymbol{\mu}_{k+1}^{i}$ $=$ $\boldsymbol{\mu}_{k}^{i}+\boldsymbol{z}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{i},$ (9c)
$\boldsymbol{\nu}_{k+1}^{i}$ $=$ $\boldsymbol{\nu}_{k}^{i}+\frac{1}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left(\boldsymbol{x}_{k+1}^{i}-\boldsymbol{x}_{k+1}^{j}\right),$ (9d)
$\boldsymbol{\xi}_{k+1}^{i}$ $=$ $\boldsymbol{\xi}_{k}^{i}+\frac{1}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left(\boldsymbol{w}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{j}\right).$ (9e)
where the local constraint set $\Sigma^{i}$, for $i\in\mathcal{V}$ as
$\Sigma^{i}=\left\{(\boldsymbol{x},\boldsymbol{w})\in\mathcal{X}\times[0,1]^{m}:g^{i}(\boldsymbol{x},\boldsymbol{w})\leq\boldsymbol{0}_{q}\right\},$ (10)
and $N^{-}(i)$ and $N^{+}(i)$ denote the sets of in-neighbors and out-neighbors of node $i\in\mathcal{V}$, respectively.
For the derivation of the above algorithm see Appendix Appendix: Derivation of the Distributed ADMM. Note that problem (9) is a convex optimization problem that can be handled with optimization solvers. As for the centralized case, problem (9b) is equivalent to solving an MWRA problem. We have
Proposition II.2
For all agents $i\in\mathcal{V}$, problem (9b) is equivalent to solving the following discrete optimization problem at each iteration $k\geq 0$:
$\boldsymbol{z}^{i}_{k+1}=\underset{\boldsymbol{z}\in\mathcal{Z}}{\arg\min}\;\boldsymbol{z}^{T}\boldsymbol{h}^{i}_{k},$ (11)
where the weight vector $\boldsymbol{h}^{i}_{k}\in\mathbb{R}^{m}$ is given by
$\boldsymbol{h}_{k}^{i}=\boldsymbol{\mu}_{k}^{i}-\boldsymbol{w}_{k+1}^{i},\text{ }i\in\mathcal{V}.$
Proof:
The proof is similar to Proof II-B. ∎
Remark II.3
When the graph is undirected, the above algorithm still works by observing that $N^{-}(i)=N^{+}(i)$, for all $i\in\mathcal{V}$. It takes the following form:
$(\boldsymbol{x}^{i},\boldsymbol{w}^{i})_{k+1}=\underset{(\boldsymbol{x}^{i},\boldsymbol{w}^{i})\in\Sigma^{i}}{\arg\min}\left\{f^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i})\right$.
$+\frac{\rho}{2}\left\|\boldsymbol{z}_{k}^{i}-\boldsymbol{w}^{i}+\boldsymbol{\mu}_{k}^{i}\right\|^{2}+(\boldsymbol{\nu}_{k}^{i})^{\top}\boldsymbol{x}^{i}$
$+(\boldsymbol{\xi}_{k}^{i})^{\top}\boldsymbol{w}^{i}+\rho\sum_{j\in N(i)}\left\|\boldsymbol{x}^{i}-\frac{\boldsymbol{x}_{k}^{i}+\boldsymbol{x}_{k}^{j}}{2}\right\|^{2}$
$+\left.\rho\sum_{j\in N(i)}\left\|\boldsymbol{w}^{i}-\frac{\boldsymbol{w}_{k}^{i}+\boldsymbol{w}_{k}^{j}}{2}\right\|^{2}\right\},$
$\boldsymbol{z}_{k+1}^{i}$ $=$ $\underset{\boldsymbol{z}^{i}\in\mathcal{Z}}{\arg\min}\left\|\boldsymbol{z}^{i}-\boldsymbol{w}_{k+1}^{i}+\boldsymbol{\mu}_{k}^{i}\right\|^{2},$ (12b)
$\boldsymbol{\mu}_{k+1}^{i}$ $=$ $\boldsymbol{\mu}_{k}^{i}+\boldsymbol{z}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{i},$ (12c)
$\boldsymbol{\nu}_{k+1}^{i}$ $=$ $\boldsymbol{\nu}_{k}^{i}+\sum_{j\in N(i)}\left(\boldsymbol{x}_{k+1}^{i}-\boldsymbol{x}_{k+1}^{j}\right),$ (12d)
$\boldsymbol{\xi}_{k+1}^{i}$ $=$ $\boldsymbol{\xi}_{k}^{i}+\sum_{j\in N(i)}\left(\boldsymbol{w}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{j}\right).$ (12e)
II-D Execution Details
Initialization.
As noted in [17, 23], ADMM on nonconvex models is sensitive to initialization and to the penalty parameter $\rho$. We adopt a relaxed start by sampling $\mathbf{w}_{0}$ from the convex hull of $\mathcal{Z}$. In the distributed case, we initialize all agents identically (e.g., $\boldsymbol{w}_{0}^{\,i}=\bar{\boldsymbol{w}}$, $\boldsymbol{x}_{0}^{\,i}=\bar{\boldsymbol{x}}$ for all $i$) to reduce transient disagreement and to speed up the convergence. Nonetheless, convergence and consensus ultimately depend on the update rules and graph connectivity [37]. The penalty parameter $\rho$ trades off feasibility and optimality: a larger $\rho$ enforces constraints more strictly, while a smaller $\rho$ favors objective improvement.
Computational Cost.
In the centralized setting, in each iteration of ADMM, the main computational burden lies in solving the convex subproblem (5a). By contrast, the projection step (5b) onto the tree constraint is significantly more efficient, which can be solved exactly in a polynomial time in both directed and undirected graphs. In the distributed setting, the same complexity occurs for each agent $i\in\mathcal{V}$.
Convergence.
Because $\mathcal{Z}$ is a nonconvex constraint set, global convergence of ADMM is not guaranteed in general [17]. In practice, however, we observe that the method yields high-quality approximate solutions for our application.
To monitor progress in the centralized setting, we define the total residual at iteration $k\geq 1$ as
$\bigl{\|}\boldsymbol{\mu}_{k}-\boldsymbol{\mu}_{k-1}\bigr{\|}\;+\;\bigl{\|}(\boldsymbol{x},\boldsymbol{w})_{k}-(\boldsymbol{x},\boldsymbol{w})_{k-1}\bigr{\|},$ (13)
which aggregates dual and primal changes. Algorithm (5a)–(5c) is declared converged when it is below the tolerance parameter.
In the distributed setting, we define the residual
$\frac{1}{n}\sum_{i\in\mathcal{V}}\left\|(\boldsymbol{\mu}^{i},\boldsymbol{\nu}^{i},\boldsymbol{\xi}^{i})_{k}-(\boldsymbol{\mu}^{i},\boldsymbol{\nu}^{i},\boldsymbol{\xi}^{i})_{k-1}\right\|$ (14)
$+\frac{1}{n}\sum_{i\in\mathcal{V}}\left\|(\boldsymbol{x}^{i},\boldsymbol{w}^{i})_{k}-(\boldsymbol{x}^{i},\boldsymbol{w}^{i})_{k-1}\right\|,$
and stop when the ”average” residual (14) is below the tolerance error parameter.
Remark II.4
For clarity of presentation, we adopt the centralized stopping criterion (14), which contrasts with the inherently distributed nature of Algorithm (9)–(9e). The choice is motivated by the fact that a fully distributed stopping rule may lead each agent to terminate at different iterations, making the results less straightforward to interpret numerically. Nevertheless, the proposed framework can be adapted to incorporate a distributed stopping criterion, as discussed in [38].
III Application: Multicommodity Flow Formulations for Spanning Trees with Hop Constraints
In this section, we execute both ADMM-based algorithms, the centralized (5a)–(5c) and the distributed (9)–(9e), and compare their quality with the solutions obtained using the commercial solver Gurobi.
All numerical experiments were performed on a workstation equipped with an Intel® Core™ i7-8650U CPU (4 cores, 8 threads, base frequency 1.90  GHz, turbo boost up to 4.20  GHz) and 16  GB of DDR4 RAM.
To measure the deviation of the ADMM solution from the optimal one, we define the optimality gap as
$\textrm{Gap}=\left(\frac{\textrm{ADMM Obj}}{\textrm{Gurobi Obj}}-1\right)\times 100\%,$ (15)
where ADMM Obj and Gurobi Obj denote the objective values obtained by the ADMM algorithms and Gurobi, respectively. A positive gap indicates that the ADMM result is suboptimal relative to Gurobi, while $\textrm{Gap}=0$ denotes an exact match.
III-A The Model
Let $G=(\mathcal{V},\mathcal{E})$ be an undirected graph with $|\mathcal{E}|=m$, where each edge $(i,j)\in\mathcal{E}$ has a nonnegative cost $c_{ij}\in\mathbb{R}_{\geq 0}$. For flow modeling, we work with the bidirected arc set
$\mathcal{A}\;=\;\{(i,j),\,(j,i)\,:\,(i,j)\in\mathcal{E}\}.$
We define a set of commodities $F$; each commodity $f\in F$ has an origin $O(f)\in\mathcal{V}$ and a destination $D(f)\in\mathcal{V}$. The goal is to find a minimum-cost spanning tree (on $\mathcal{E}$) that supports routing all commodities (over $\mathcal{A}$) while ensuring that each commodity’s path length does not exceed a bound $d\in\mathbb{N}$ (see e.g., [1, 2, 3, 4]).
The binary variable $z_{ij}\in\{0,1\}$ indicates whether undirected edge $(i,j)\in\mathcal{E}$ is in the tree (we write $z_{ij}=z_{ji}$ for $(i,j)\in\mathcal{E}$). For each commodity $f\in F$, the arc-flow variables $y_{ij}^{f}\in\{0,1\}$ indicate whether commodity $f$ uses arc $(i,j)\in\mathcal{A}$. We let $\boldsymbol{y}^{f}\in\{0,1\}^{2m}$ collect all flows for $f$. The model reads [1, 2, 3]:
$\underset{(\boldsymbol{y},\boldsymbol{z})}{\text{minimize}}\sum_{(i,j)\in\mathcal{E}}c_{ij}\,z_{ij},$ (16)
subject to
$\sum_{(i,j)\in\mathcal{E}}z_{ij}=n-1,\text{ }\sum_{\mathcal{E}(S)}z_{ij}\leq|S|-1,\ S\subset\mathcal{V},\ |S|\geq 2,$ (17a)
$\sum_{j\in N^{-}(i)}y_{ji}^{f}-\sum_{j\in N^{+}(i)}y_{ij}^{f}\;$ (17b)
$=\;\begin{cases}1&\text{if }i=O(f),\\ 0&\text{if }i\notin\{O(f),D(f)\},\\ -1&\text{if }i=D(f),\end{cases}\quad\forall i\in\mathcal{V},\ \forall f\in F,$
$y_{ij}^{f}+y_{ji}^{f}\;\leq\;z_{ij},\quad\forall(i,j)\in\mathcal{E},\ \forall f\in F,$ (17c)
$\sum_{(i,j)\in\mathcal{A}}y_{ij}^{f}\;\leq\;d,\quad\forall f\in F,$ (17d)
$y_{ij}^{f}\in\{0,1\},\ \forall(i,j)\in\mathcal{A},\ \forall f\in F,$ (17e)
$\qquad z_{ij}$ $\in\{0,1\},\ \forall(i,j)\in\mathcal{E},$
where $\mathcal{E}(S)$ denotes the set of edges whose endpoints both lie in $S$. Constraint (17a) ensures that the selected edges $\boldsymbol{z}$ form a connected spanning tree over $\mathcal{E}$; see, for example, [39, 40].
III-B The Centralized Algorithm
Next, we define a relaxed version of the model, where $w_{ij}\in[0,1]$ and $u_{ij}^{f}\in[0,1]$ replace $z_{ij}$ and $y_{ij}^{f}$, respectively:
$\underset{(\boldsymbol{u},\boldsymbol{w})}{\text{minimize}}\sum_{(i,j)\in\mathcal{E}}c_{ij}\,w_{ij},$
subject to
$\sum_{(i,j)\in\mathcal{E}}w_{ij}=n-1,\text{ }\sum_{\mathcal{E}(S)}w_{ij}\leq|S|-1,\ S\subset\mathcal{V},\ |S|\geq 2,$ (18a)
$\boldsymbol{w}=\boldsymbol{z},\quad\boldsymbol{u}^{f}=\boldsymbol{y}^{f},$ (18b)
$\sum_{j\in N^{-}(i)}u_{ji}^{f}-\sum_{j\in N^{+}(i)}u_{ij}^{f}$ (18c)
$=\begin{cases}1&\text{if }i=O(f),\\[1.99997pt] 0&\text{if }i\notin\{O(f),D(f)\},\\[1.99997pt] -1&\text{if }i=D(f),\end{cases}\quad\forall i\in\mathcal{V},\ \forall f\in F,$
$u_{ij}^{f}+u_{ji}^{f}\leq w_{ij},\quad\forall(i,j)\in\mathcal{E},\ \forall f\in F,$ (18d)
$\sum_{(i,j)\in\mathcal{A}}u_{ij}^{f}\leq d,\quad\forall f\in F,$ (18e)
$w_{ij}\in[0,1],\quad\forall(i,j)\in\mathcal{E},$
$u_{ij}^{f}\in[0,1],\quad\forall(i,j)\in\mathcal{E},\ \forall f\in F.$ (18f)
The ADMM updates are defined as follows:
$(\boldsymbol{u},\boldsymbol{w})_{k+1}$ $=\arg\min_{(\boldsymbol{u},\boldsymbol{w})\in\Sigma}\left\{\sum_{(i,j)\in\mathcal{E}}c_{ij}\,w_{ij}+\tfrac{\rho}{2}\,\|\boldsymbol{z}_{k}-\boldsymbol{w}+\boldsymbol{\mu}_{k}\|^{2}\right$. (19a)
$+\left.\tfrac{\rho}{2}\,\|\boldsymbol{y}_{k}-\boldsymbol{u}+\boldsymbol{\eta}_{k}\|^{2},\right\}$
$\boldsymbol{z}_{k+1}$ $=\arg\min_{\boldsymbol{z}\in\mathcal{Z}}\ \|\boldsymbol{z}-\boldsymbol{w}_{k+1}+\boldsymbol{\mu}_{k}\|^{2},$ (19b)
$\boldsymbol{y}_{k+1}$ $=\arg\min_{\boldsymbol{y}\in\{0,1\}^{2m|F|}}\ \|\boldsymbol{y}-\boldsymbol{u}_{k+1}+\boldsymbol{\eta}_{k}\|^{2},$ (19c)
$\boldsymbol{\mu}_{k+1}$ $=\boldsymbol{\mu}_{k}+\boldsymbol{z}_{k+1}-\boldsymbol{w}_{k+1},$ (19d)
$\boldsymbol{\eta}_{k+1}$ $=\boldsymbol{\eta}_{k}+\boldsymbol{y}_{k+1}-\boldsymbol{u}_{k+1}.$ (19e)
We define the constraint set $\Sigma$ by:
[…]
Note that constraint (18a) is excluded from the set $\Sigma$ because, by Proposition II.1, problem (19b) reduces to an MST problem with weights defined in (6) which guarantees that $\boldsymbol{z}_{k}$ induces a tree for all $k\geq 1$. The second integer problem (19c) is simply a straightforward component-wise projection of $\boldsymbol{u}_{k+1}-\boldsymbol{\eta}_{k}$ onto $\{0,1\}^{2m|F|}$.
III-C The Distributed Algorithm
To decentralize problem (16)–(17e), we express the objective function and constraints as:
$\sum_{(i,j)\in\mathcal{E}}c_{ij}w_{ij}=\frac{1}{2}\sum_{i\in\mathcal{V}}\left(\sum_{j\in N(i)}c_{ij}w_{ij}\right)=:\sum_{i\in\mathcal{V}}f^{i}(\boldsymbol{x},\boldsymbol{w}).$ (20)
For the constraints in (18b), we express them in terms of local variables:
$\boldsymbol{w}^{i}$ $=$ $\boldsymbol{z}^{i},\quad(\boldsymbol{u}^{f})^{i}=(\boldsymbol{y}^{f})^{i},\quad i\in\mathcal{V},\quad$
$\boldsymbol{w}^{i}$ $=$ $\boldsymbol{w}^{j},\quad(\boldsymbol{u}^{f})^{i}=(\boldsymbol{u}^{f})^{j},\quad f\in F,\ (i,j)\in\mathcal{E}.$
The constraint (18c) can be decentralized for each agent $i\in\mathcal{V}$ as:
$\sum_{j\in N^{-}(i)}\left(u_{ji}^{f}\right)^{i}-\sum_{j\in N^{+}(i)}\left(u_{ij}^{f}\right)^{i}$
$=$ $\begin{cases}1&\text{if }i=O(f),\\ 0&\text{if }i\notin\{O(f),D(f)\},\\ -1&\text{if }i=D(f),\end{cases}\quad\forall i\in\mathcal{V},\ \forall f\in F.$
For constraint (18d), we decentralize it as:
$\left(u_{ij}^{f}\right)^{i}+\left(u_{ji}^{f}\right)^{i}\leq\left(w_{ij}\right)^{i},\quad\forall j\in N(i),\ \forall i\in\mathcal{V},\ \forall f\in F.$ (22)
For the diameter constraint (18f), we assume that for each agent $l\in\mathcal{V}$, the following local constraint is satisfied:
$\sum_{(i,j)\in\mathcal{A}}\left(u_{ij}^{f}\right)^{l}\leq d,\quad\forall f\in F,\ l\in\mathcal{V}.$ (23)
Thus, the local constraint set for each agent $i\in\mathcal{V}$ is defined as:
[…]
By adapting the distributed algorithm for this problem, we get the following ADMM updates:
$(\boldsymbol{u}^{i},\boldsymbol{w}^{i})_{k+1}$
$=\underset{(\boldsymbol{w}^{i},\boldsymbol{u}^{i})\in\Sigma^{i}}{\arg\min}\Bigg{\{}f^{i}(\boldsymbol{w}^{i})+\frac{\rho}{2}\left\|\boldsymbol{z}_{k}^{i}-\boldsymbol{w}^{i}+\boldsymbol{\mu}_{k}^{i}\right\|^{2}$
$+\frac{\rho}{2}\left\|\boldsymbol{y}_{k}^{i}-\boldsymbol{u}^{i}+\boldsymbol{\eta}_{k}^{i}\right\|^{2}+(\boldsymbol{\nu}_{k}^{i})^{\top}\boldsymbol{u}^{i}$
$+\rho\sum_{j\in N(i)}\left\|\boldsymbol{u}^{i}-\frac{\boldsymbol{u}_{k}^{i}+\boldsymbol{u}_{k}^{j}}{2}\right\|^{2}+(\boldsymbol{\xi}_{k}^{i})^{\top}\boldsymbol{w}^{i}$
$+\rho\sum_{j\in N(i)}\left\|\boldsymbol{w}^{i}-\frac{\boldsymbol{w}_{k}^{i}+\boldsymbol{w}_{k}^{j}}{2}\right\|^{2}\Bigg{\}},$
$\boldsymbol{z}_{k+1}^{i}$ $=\underset{\boldsymbol{z}^{i}\in\mathcal{Z}}{\arg\min}\left\|\boldsymbol{z}^{i}-\boldsymbol{w}_{k+1}^{i}+\boldsymbol{\mu}_{k}^{i}\right\|^{2},$
$\boldsymbol{y}_{k+1}^{i}$ $=\arg\min_{\boldsymbol{y}^{i}\in\{0,1\}^{2m|F|}}\left\|\boldsymbol{y}^{i}-\boldsymbol{u}_{k+1}^{i}+\boldsymbol{\eta}_{k}^{i}\right\|^{2},$
$\boldsymbol{\nu}_{k+1}^{i}$ $=\boldsymbol{\nu}_{k}^{i}+\sum_{j\in N(i)}\left(\boldsymbol{u}_{k+1}^{i}-\boldsymbol{u}_{k+1}^{j}\right),$
$\boldsymbol{\xi}_{k+1}^{i}$ $=\boldsymbol{\xi}_{k}^{i}+\sum_{j\in N(i)}\left(\boldsymbol{w}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{j}\right),$
$\boldsymbol{\mu}_{k+1}^{i}$ $=\boldsymbol{\mu}_{k}^{i}+\boldsymbol{z}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{i},$
$\boldsymbol{\eta}_{k+1}^{i}$ $=\boldsymbol{\eta}_{k}^{i}+\boldsymbol{y}_{k+1}^{i}-\boldsymbol{u}_{k+1}^{i}.$
III-D The Numerical Results
We evaluate the performances of the centralized algorithm (5a)–(5c) and the distributed one (9)–(9e) on Erdős–Rényi random graphs [41] with connectivity ratio $p=0.5$, using network sizes $n\in\{10,50,100\}$, and $|F|=\left\lfloor\frac{n}{5}\right\rfloor$ randomly sampled commodities. The diameter $d$ is selected so that the problem is feasible. The initial point is $\boldsymbol{w}_{0}=\boldsymbol{w}^{i}_{0}=\mathbf{1}_{m}$, for all $i\in\mathcal{V}$. The ADMM solution quality is evaluated against the global optimum from Gurobi [42]. The error tolerance parameter is set to $10^{-4}$.
The simulations compare three key aspects: the number of iterations required to reach convergence (Figures 13), the objective value evolution (Figures 46), and the agents’ trajectories in the distributed setting (Figures 79). For each graph size $n$, the algorithms are executed three times using different penalty parameters, $\rho\in\{0.1,1,10\}$.
(a) $n=10$, $\rho=0.1$
(b) $n=10$, $\rho=1$
(c) $n=10$, $\rho=10$
Figure 1: Relative residual errors defined in (13) and (14) for centralized and distributed ADMM with $n=10$ across $\rho\in\{0.1,1,10\}$.
(a) $n=50$, $\rho=0.1$
(b) $n=50$, $\rho=1$
(c) $n=50$, $\rho=10$
Figure 2: Relative residual errors defined in (13) and (14) for centralized and distributed ADMM with $n=50$ across $\rho\in\{0.1,1,10\}$.
(a) $n=100$, $\rho=0.1$
(b) $n=100$, $\rho=1$
(c) $n=100$, $\rho=10$
Figure 3: Relative residual errors defined in (13) and (14) for centralized and distributed ADMM with $n=100$ across $\rho\in\{0.1,1,10\}$.
Residual Error
As shown in Figures 13, for both centralized and distributed ADMM, the relative residual errors decrease monotonically in most runs, with convergence speed governed by the penalty parameter $\rho$ and network size $n$. Small penalties ($\rho=0.1$) give smooth but slow decay with long tails; moderate penalties ($\rho=1$) achieve rapid, stable convergence; large penalties ($\rho=10$) produce the fastest initial drop but can introduce mild oscillations, especially in distributed runs. Increasing $n$ slows distributed convergence due to longer consensus times, while centralized performance is largely unaffected. Across all settings, centralized residuals fall faster, but after sufficient iterations, both approaches reach a similar residual floor, illustrating the trade-off between speed and communication overhead; moderate $\rho$ values generally balance stability and speed best.
(a) $n=10$, $\rho=0.1$
(b) $n=10$, $\rho=1$
(c) $n=10$, $\rho=10$
Figure 4: Objective $\boldsymbol{c}^{\top}\boldsymbol{w}_{k}$ (distributed vs. centralized) for $n=10$ across $\rho\in\{0.1,1,10\}$.
(a) $n=50$, $\rho=0.1$
(b) $n=50$, $\rho=1$
(c) $n=50$, $\rho=10$
Figure 5: Objective $\boldsymbol{c}^{\top}\boldsymbol{w}_{k}$ (distributed vs. centralized) for $n=50$ and $\rho\in\{0.1,1,10\}$.
(a) $n=100$, $\rho=0.1$
(b) $n=100$, $\rho=1$
(c) $n=100$, $\rho=10$
Figure 6: Objective $\boldsymbol{c}^{\top}\boldsymbol{w}_{k}$ (distributed vs. centralized) for $n=100$ across $\rho\in\{0.1,1,10\}$.
Objective Evolution
Based on Figures 46, the objective value $\boldsymbol{c}^{\top}\boldsymbol{w}_{k}$ for both centralized and distributed ADMM generally exhibits similar trends, with the distributed curve tracking the centralized one after a short initial lag. For fixed $n$, a small penalty ($\rho=0.1$) produces a slow, mostly monotonic decrease, eventually aligning with the centralized trajectory; a moderate penalty ($\rho=1$) accelerates convergence and reduces oscillations, enabling earlier alignment; a large penalty ($\rho=10$) yields the fastest initial drop but results in a persistent gap between the two methods. This behavior suggests that, in the high-penalty regime, the centralized algorithm tends to prioritize feasibility enforcement over continued objective minimization, whereas the distributed method continues to seek further improvement toward optimality even when $\rho$ is large. For fixed $\rho$, increasing $n$ from $10$ to $50$ to $100$ prolongs the alignment lag due to slower information propagation in larger communication graphs. When $\rho\in\{0.1,1\}$, the distributed objective ultimately aligns with the centralized one despite the delay, whereas for $\rho=10$ the gap remains for all tested sizes. Overall, centralized updates stabilize in fewer iterations, while distributed updates exhibit a short tracking delay before alignment except in the high-penalty case, where the gap gets larger between the optimal values.
(a) $n=10$, $\rho=0.1$
(b) $n=10$, $\rho=1$
(c) $n=10$, $\rho=10$
Figure 7: Per-agent norms $\|(\boldsymbol{w}^{i},\boldsymbol{u}^{i})_{k}\|^{2}$ for $n=10$ across $\rho\in\{0.1,1,10\}$.
(a) $n=50$, $\rho=0.1$
(b) $n=50$, $\rho=1$
(c) $n=50$, $\rho=10$
Figure 8: Per-agent norms $\|(\boldsymbol{w}^{i},\boldsymbol{u}^{i})_{k}\|^{2}$ for $n=50$ and $\rho\in\{0.1,1,10\}$.
(a) $n=100$, $\rho=0.1$
(b) $n=100$, $\rho=1$
(c) $n=100$, $\rho=10$
Figure 9: Per-agent norms $\|(\boldsymbol{w}^{i},\boldsymbol{u}^{i})_{k}\|^{2}$ for $n=100$ across $\rho\in\{0.1,1,10\}$.
Agent’s Behavior
As shown in Figures 79, across all experiments, the per–agent trajectories $\|(\boldsymbol{w}^{i},\boldsymbol{u}^{i})_{k}\|^{2}$ quickly cluster and then evolve in near unison, indicating effective consensus in the distributed scheme. For fixed $n$, a small penalty ($\rho=0.1$) leads to slow but smooth synchronization with little overshoot; a moderate penalty ($\rho=1$) achieves faster clustering and stable alignment; a large penalty ($\rho=10$) gives the quickest transient and earliest clustering, sometimes with mild oscillations. For fixed $\rho$, increasing $n$ from $10$ to $50$ to $100$ extends clustering time due to slower information spread, yet agents still converge to nearly identical trajectories after a short transient.
ADMM vs. Gurobi
This section evaluates the execution time and solution quality of centralized and distributed ADMM against the exact solutions obtained by Gurobi, across different problem sizes $n$ and penalty parameters $\rho$, as summarized in Tables IIII.
TABLE I: ADMM vs Gurobi for varying $\rho$ on network with $n=10$.
$\rho$ Centr. ADMM Dist. ADMM Time Gurobi
Gap (%) Time Gap (%) Time
0.1 1.32% 4.76 1.32% 11.23 4.58
1 1.33% 2.76 1.32% 9.24 4.58
10 15.57% 2.83 5.32% 6.24 4.58
TABLE II: ADMM vs Gurobi for varying $\rho$ on network with $n=50$.
$\rho$ Cent. ADMM Dist. ADMM Time Gurobi
Gap (%) Time Gap (%) Time
0.1 4.57% 11.65 4.57% 21.01 8.34
1 4.57% 9.11 4.57% 25.23 8.34
10 49.32% 6.87 15.79% 14.92 8.34
TABLE III: ADMM vs Gurobi for varying $\rho$ on network with $n=100$.
$\rho$ Cent. ADMM Dist. ADMM Time Gurobi
Gap (%) Time Gap (%) Time
0.1 8.53% 23.43 8.53% 103.4 16.98
1 8.83% 17.60 8.53% 72.11 16.98
10 131.17% 9.54 26.76% 47.21 16.98
[…]
Distributed ADMM: runtimes are longer due to consensus overhead (e.g., $103.4$ s for $n=100$ at $\rho=0.1$), while the gaps for small or moderate penalties remain close to those of the centralized case. Notably, for $\rho=10$, the distributed method is more robust: gaps are $5.32\%$ at $n=10$, $15.79\%$ at $n=50$, and $26.76\%$ at $n=100$, compared to much larger gaps in the centralized runs. This indicates that, under large penalty values, the distributed scheme continues to improve objective quality despite slower convergence, whereas the centralized method tends to prioritize feasibility enforcement.
Conclusion
We proposed both centralized and distributed ADMM-based heuristics to solve mixed-integer network design problems with spanning tree or rooted arborescence constraints. Feasibility is guaranteed at each iteration by projecting onto the exact MST or MWRA solution. When applied to hop-constrained multicommodity flow problems, the methods produced near-optimal solutions with small optimality gaps and achieved substantial runtime savings compared to Gurobi. The centralized approach converged more quickly, while the distributed version showed greater robustness under large penalty parameters. Overall, the framework is scalable and reliable for large-scale network optimization, with potential extensions including adaptive penalty tuning, asynchronous communication, and support for additional graph constraints.
Appendix: Derivation of the Distributed ADMM
First, we derive the distributed version of the ADMM algorithm on directed graphs. Then, we deduce the algorithm on undirected graphs by considering both directions on the edges.
Introduce the variables $\boldsymbol{t}^{ij},\boldsymbol{s}^{ij}\in\mathbb{R}^{m}$, where $(i,j)\in\mathcal{A}$. Problem (8) is equivalent to:
$\begin{array}[]{ll}\text{minimize}&\sum_{i\in\mathcal{V}}f^{i}(\boldsymbol{x}^{i},\boldsymbol{z}^{i})\\ \text{subject to}&\sum_{i\in\mathcal{V}}g^{i}(\boldsymbol{x}^{i},\boldsymbol{z}^{i})\leq\boldsymbol{0}_{q},\\ &\boldsymbol{w}^{i}=\boldsymbol{z}^{i},\\ &\boldsymbol{t}^{ji}=\boldsymbol{t}^{ij}=\boldsymbol{x}^{i},\quad\boldsymbol{s}^{ji}=\boldsymbol{s}^{ij}=\boldsymbol{w}^{i},\end{array}$ (25)
and consider the Lagrangian:
$\mathcal{L}_{\rho}(\boldsymbol{x},\boldsymbol{w},\boldsymbol{z},\boldsymbol{\mu},\boldsymbol{t},\boldsymbol{s},\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$
$=$ $\sum_{i\in\mathcal{V}}f^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i})+\sum_{i\in\mathcal{V}}\boldsymbol{\mu}^{i\top}(\boldsymbol{z}^{i}-\boldsymbol{w}^{i})$
$+\frac{\rho}{2}\sum_{i\in\mathcal{V}}\left\|\boldsymbol{z}^{i}-\boldsymbol{w}^{i}\right\|^{2}+\sum_{(i,j)\in\mathcal{A}}(\boldsymbol{\alpha}^{ij})^{\top}(\boldsymbol{x}^{i}-\boldsymbol{t}^{ij})$
$+\frac{\rho}{2}\sum_{(i,j)\in\mathcal{A}}\left\|\boldsymbol{x}^{i}-\boldsymbol{t}^{ij}\right\|^{2}+\sum_{(i,j)\in\mathcal{A}}(\boldsymbol{\beta}^{ij})^{\top}(\boldsymbol{x}^{i}-\boldsymbol{t}^{ij})$
$+\frac{\rho}{2}\sum_{(i,j)\in\mathcal{A}}\left\|\boldsymbol{x}^{i}-\boldsymbol{t}^{ji}\right\|^{2}+\sum_{(i,j)\in\mathcal{A}}(\boldsymbol{\gamma}^{ij})^{\top}(\boldsymbol{w}^{i}-\boldsymbol{s}^{ij})$
$+\frac{\rho}{2}\sum_{(i,j)\in\mathcal{A}}\left\|\boldsymbol{w}^{i}-\boldsymbol{s}^{ij}\right\|^{2}+\sum_{(i,j)\in\mathcal{A}}(\boldsymbol{\delta}^{ij})^{\top}(\boldsymbol{w}^{i}-\boldsymbol{s}^{ji})$
$+\frac{\rho}{2}\sum_{(i,j)\in\mathcal{A}}\left\|\boldsymbol{w}^{i}-\boldsymbol{s}^{ji}\right\|^{2}.$
The Lagrangian can be written in separable form as:
$\mathcal{L}_{\rho}(\boldsymbol{x},\boldsymbol{w},\boldsymbol{z},\boldsymbol{\mu},\boldsymbol{t},\boldsymbol{s},\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$
$=$ $\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{N}^{+}(i)}\mathcal{L}_{\rho}^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i},\boldsymbol{z}^{i},\boldsymbol{\mu}^{i},\boldsymbol{t}^{ij},\boldsymbol{s}^{ij},\boldsymbol{\alpha}^{ij},\boldsymbol{\beta}^{ij},\boldsymbol{\gamma}^{ij},\boldsymbol{\delta}^{ij})$
$=$ $\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{N}^{-}(i)}\mathcal{L}_{\rho}^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i},\boldsymbol{z}^{i},\boldsymbol{\mu}^{i},\boldsymbol{t}^{ij},\boldsymbol{s}^{ij},\boldsymbol{\alpha}^{ij},\boldsymbol{\beta}^{ij},\boldsymbol{\gamma}^{ij},\boldsymbol{\delta}^{ij}).$
where $\mathcal{L}_{\rho}^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i},\boldsymbol{z}^{i},\boldsymbol{\mu}^{i},\boldsymbol{t}^{ij},\boldsymbol{s}^{ij},\boldsymbol{\alpha}^{ij},\boldsymbol{\beta}^{ij},\boldsymbol{\gamma}^{ij},\boldsymbol{\delta}^{ij})$ is given by:
$\mathcal{L}_{\rho}^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i},\boldsymbol{z}^{i},\boldsymbol{\mu}^{i},\boldsymbol{t}^{ij},\boldsymbol{s}^{ij},\boldsymbol{\alpha}^{ij},\boldsymbol{\beta}^{ij},\boldsymbol{\gamma}^{ij},\boldsymbol{\delta}^{ij})$
$=$ $f^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i})+\boldsymbol{\mu}^{i\top}(\boldsymbol{z}^{i}-\boldsymbol{w}^{i})+\frac{\rho}{2}\left\|\boldsymbol{z}^{i}-\boldsymbol{w}^{i}\right\|^{2}$
$+\sum_{j\in\mathcal{N}^{+}(i)}(\boldsymbol{\alpha}^{ij})^{\top}(\boldsymbol{x}^{i}-\boldsymbol{t}^{ij})+\frac{\rho}{2}\sum_{j\in\mathcal{N}^{+}(i)}\left\|\boldsymbol{x}^{i}-\boldsymbol{t}^{ij}\right\|^{2}$
$+\sum_{j\in\mathcal{N}^{+}(i)}(\boldsymbol{\beta}^{ij})^{\top}(\boldsymbol{x}^{j}-\boldsymbol{t}^{ij})+\frac{\rho}{2}\sum_{j\in\mathcal{N}^{+}(i)}\left\|\boldsymbol{x}^{j}-\boldsymbol{t}^{ij}\right\|^{2}$
$+\sum_{j\in\mathcal{N}^{+}(i)}(\boldsymbol{\gamma}^{ij})^{\top}(\boldsymbol{w}^{i}-\boldsymbol{s}^{ij})+\frac{\rho}{2}\sum_{j\in\mathcal{N}^{+}(i)}\left\|\boldsymbol{w}^{i}-\boldsymbol{s}^{ij}\right\|^{2}$
$+\sum_{j\in\mathcal{N}^{+}(i)}(\boldsymbol{\delta}^{ij})^{\top}(\boldsymbol{w}^{j}-\boldsymbol{s}^{ij})+\frac{\rho}{2}\sum_{j\in\mathcal{N}^{+}(i)}\left\|\boldsymbol{w}^{i}-\boldsymbol{s}^{ij}\right\|^{2}.$
The ADMM iterations for all $i\in\mathcal{V}$ and $j\in\mathcal{N}^{+}(i)$ are given by:
$(\boldsymbol{x}^{i},\boldsymbol{w}^{i})_{k+1}$ $=$ $\underset{(\boldsymbol{x}^{i},\boldsymbol{w}^{i})\in\Sigma^{i}}{\arg\min}\mathcal{L}_{\rho}^{i}\left(\begin{array}[]{c}\boldsymbol{x}^{i},\boldsymbol{w}^{i},\boldsymbol{z}_{k}^{i},\boldsymbol{\mu}_{k}^{i},\boldsymbol{t}_{k}^{ij},\\ \boldsymbol{s}_{k}^{ij},\boldsymbol{\alpha}_{k}^{ij},\boldsymbol{\beta}_{k}^{ij},\boldsymbol{\gamma}_{k}^{ij},\boldsymbol{\delta}_{k}^{ij}\end{array}\right),$
$(\boldsymbol{s}^{ij},\boldsymbol{t}^{ij})_{k+1}$ $=$ $\arg\min\mathcal{L}_{\rho}^{i}\left(\begin{array}[]{c}\boldsymbol{x}_{k+1}^{i},\boldsymbol{w}_{k+1}^{i},\boldsymbol{z}_{k}^{i},\boldsymbol{\mu}_{k}^{i},\boldsymbol{t}^{ij},\\ \boldsymbol{s}^{ij},\boldsymbol{\alpha}_{k}^{ij},\boldsymbol{\beta}_{k}^{ij},\boldsymbol{\gamma}_{k}^{ij},\boldsymbol{\delta}_{k}^{ij}\end{array}\right),$
$\boldsymbol{z}_{k+1}^{i}$ $=$ $\underset{\boldsymbol{z}^{i}\in\mathcal{Z}}{\arg\min}\left\|\boldsymbol{z}^{i}-\boldsymbol{w}_{k+1}^{i}+\boldsymbol{\mu}_{k}^{i}\right\|_{\mathbb{R}^{m}}^{2},$
$\boldsymbol{\mu}_{k+1}^{i}$ $=$ $\boldsymbol{\mu}_{k}^{i}+\rho\left(\boldsymbol{z}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{i}\right),$
$\boldsymbol{\alpha}_{k+1}^{ij}$ $=$ $\boldsymbol{\alpha}_{k}^{ij}+\rho\left(\boldsymbol{x}_{k+1}^{i}-\boldsymbol{t}_{k+1}^{ij}\right),$
$\boldsymbol{\beta}_{k+1}^{ij}$ $=$ $\boldsymbol{\beta}_{k}^{ij}+\rho\left(\boldsymbol{x}_{k+1}^{j}-\boldsymbol{t}_{k+1}^{ij}\right),$
$\boldsymbol{\gamma}_{k+1}^{ij}$ $=$ $\boldsymbol{\gamma}_{k}^{ij}+\rho\left(\boldsymbol{w}_{k+1}^{i}-\boldsymbol{s}_{k+1}^{ij}\right),$
$\boldsymbol{\delta}_{k+1}^{ij}$ $=$ $\boldsymbol{\delta}_{k}^{ij}+\rho\left(\boldsymbol{w}_{k+1}^{j}-\boldsymbol{s}_{k+1}^{ij}\right).$
where the local constraints set $\Sigma^{i}$ is defined in (10). Since the problem is unconstrained and the objective functions are convex, and the variables $\boldsymbol{s}^{ij}$ and $\boldsymbol{t}^{ij}$ are uncoupled, we can simply take the gradient to solve for the variables:
$\nabla_{\boldsymbol{t}_{k+1}^{ij}}\mathcal{L}_{\rho}^{i}(\boldsymbol{x}_{k+1}^{i},\boldsymbol{w}_{k+1}^{i},\boldsymbol{z}_{k+1}^{i},\boldsymbol{\mu}_{k}^{i},\boldsymbol{t}^{ij},\boldsymbol{s}_{k}^{ij},\boldsymbol{\alpha}_{k}^{ij},\boldsymbol{\beta}_{k}^{ij},\boldsymbol{\gamma}_{k}^{ij},\boldsymbol{\delta}_{k}^{ij})$ $=$ $\boldsymbol{0}_{m},$
$\nabla_{\boldsymbol{s}_{k+1}^{ij}}\mathcal{L}_{\rho}^{i}(\boldsymbol{x}_{k+1}^{i},\boldsymbol{w}_{k+1}^{i},\boldsymbol{z}_{k+1}^{i},\boldsymbol{\mu}_{k}^{i},\boldsymbol{t}_{k}^{ij},\boldsymbol{s}^{ij},\boldsymbol{\alpha}_{k}^{ij},\boldsymbol{\beta}_{k}^{ij},\boldsymbol{\gamma}_{k}^{ij},\boldsymbol{\delta}_{k}^{ij})$ $=$ $\boldsymbol{0}_{m},$
which yield:
$\boldsymbol{\alpha}_{k}^{ij}-\boldsymbol{\beta}_{k}^{ji}-\rho(\boldsymbol{x}_{k+1}^{i}+\boldsymbol{x}_{k+1}^{j})+2\rho\boldsymbol{t}_{k+1}^{ij}=\boldsymbol{0}_{m},$
$\boldsymbol{\gamma}_{k}^{ij}-\boldsymbol{\delta}_{k}^{ji}-\rho(\boldsymbol{w}_{k+1}^{i}+\boldsymbol{w}_{k+1}^{j})+2\rho\boldsymbol{s}_{k+1}^{ij}=\boldsymbol{0}_{m},$
which gives:
$\boldsymbol{t}_{k+1}^{ij}$ $=$ $\frac{1}{2\rho}\boldsymbol{\alpha}_{k}^{ij}+\frac{1}{2\rho}\boldsymbol{\beta}_{k}^{ij}+\frac{1}{2}\left(\boldsymbol{x}_{k+1}^{i}+\boldsymbol{x}_{k+1}^{j}\right),$ (27a)
$\boldsymbol{s}_{k+1}^{ij}$ $=$ $\frac{1}{2\rho}\boldsymbol{\gamma}_{k}^{ij}+\frac{1}{2\rho}\boldsymbol{\delta}_{k}^{ij}+\frac{1}{2}\left(\boldsymbol{w}_{k+1}^{i}+\boldsymbol{w}_{k+1}^{j}\right).$ (27b)
From the adjoint equations, we get:
$\boldsymbol{\alpha}_{k+1}^{ij}$ $=$ $\frac{1}{2}(\boldsymbol{\alpha}_{k}^{ij}-\boldsymbol{\beta}_{k}^{ij})+\frac{\rho}{2}(\boldsymbol{x}_{k+1}^{i}-\boldsymbol{x}_{k+1}^{j}),$
$\boldsymbol{\beta}_{k+1}^{ij}$ $=$ $\frac{1}{2}(\boldsymbol{\beta}_{k}^{ij}-\boldsymbol{\alpha}_{k}^{ij})+\frac{\rho}{2}(\boldsymbol{x}_{k+1}^{j}-\boldsymbol{x}_{k+1}^{i}),$
$\boldsymbol{\gamma}_{k+1}^{ij}$ $=$ $\frac{1}{2}(\boldsymbol{\gamma}_{k}^{ij}-\boldsymbol{\delta}_{k}^{ji})+\frac{\rho}{2}(\boldsymbol{w}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{j}),$
$\boldsymbol{\delta}_{k+1}^{ij}$ $=$ $\frac{1}{2}(\boldsymbol{\delta}_{k}^{ij}-\boldsymbol{\gamma}_{k}^{ij})+\frac{\rho}{2}(\boldsymbol{w}_{k+1}^{j}-\boldsymbol{w}_{k+1}^{i}).$
Summing up yields:
$\boldsymbol{\alpha}_{k+1}^{ij}+\boldsymbol{\beta}_{k+1}^{ij}=\boldsymbol{\ 0}_{m},\quad\boldsymbol{\gamma}_{k+1}^{ij}+\boldsymbol{\delta}_{k+1}^{ij}=\boldsymbol{0}_{m},\quad\forall k\geq 0.$
Taking $\boldsymbol{\alpha}_{0}^{ij}=\boldsymbol{\beta}_{0}^{ij}=\boldsymbol{\ \gamma}_{0}^{ij}=\boldsymbol{\delta}_{0}^{ij}=\boldsymbol{0}_{q}$, we obtain:
$\boldsymbol{\alpha}_{k}^{ij}+\boldsymbol{\beta}_{k}^{ij}=\boldsymbol{0}_{q},\quad\boldsymbol{\gamma}_{k}^{ij}+\boldsymbol{\delta}_{k}^{ij}=\boldsymbol{0}_{m},\quad\forall k\geq 0.$
This yields from (27a)-(27b):
$\boldsymbol{t}_{k+1}^{ij}$ $=$ $\frac{1}{2}(\boldsymbol{x}_{k+1}^{i}+\boldsymbol{x}_{k+1}^{j}),$
$\boldsymbol{s}_{k+1}^{ij}$ $=$ $\frac{1}{2}(\boldsymbol{w}_{k+1}^{i}+\boldsymbol{w}_{k+1}^{j}).$
Thus, the dual variables become:
$\boldsymbol{\alpha}_{k+1}^{ij}$ $=$ $\boldsymbol{\alpha}_{k}^{ij}+\frac{\rho}{2}\left(\boldsymbol{x}_{k+1}^{i}-\boldsymbol{x}_{k+1}^{j}\right),$ (28a)
$\boldsymbol{\beta}_{k+1}^{ij}$ $=$ $\boldsymbol{\beta}_{k}^{ij}+\frac{\rho}{2}\left(\boldsymbol{x}_{k+1}^{j}-\boldsymbol{x}_{k+1}^{i}\right),$ (28b)
$\boldsymbol{\gamma}_{k+1}^{ij}$ $=$ $\boldsymbol{\gamma}_{k}^{ij}+\frac{\rho}{2}\left(\boldsymbol{w}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{j}\right),$ (28c)
$\boldsymbol{\delta}_{k+1}^{ij}$ $=$ $\boldsymbol{\delta}_{k}^{ij}+\frac{\rho}{2}\left(\boldsymbol{w}_{k+1}^{j}-\boldsymbol{w}_{k+1}^{i}\right).$ (28d)
To handle the terms in the local Lagrangian $\mathcal{L}_{\rho}^{i}$, we eliminate the terms depending on the dual variables from the iterate ($\boldsymbol{x}^{i},\boldsymbol{w}^{i}$).
Defining the new multipliers $\boldsymbol{\nu}_{k}^{i}$ and $\boldsymbol{\xi}_{k}^{i}$ by:
$\boldsymbol{\nu}_{k}^{i}$ $=$ $\sum_{j\in\mathcal{N}^{+}(i)}\boldsymbol{\alpha}_{k}^{ij}+\sum_{j\in\mathcal{N}^{-}(i)}\boldsymbol{\beta}_{k}^{ji},$ (29a)
$\boldsymbol{\xi}_{k}^{i}$ $=$ $\sum_{j\in\mathcal{N}^{+}(i)}\boldsymbol{\gamma}_{k}^{ij}+\sum_{j\in\mathcal{N}^{-}(i)}\boldsymbol{\delta}_{k}^{ji}.$ (29b)
By combining (28a)-(28d) and (29a)-(29b), we get:
$\boldsymbol{\nu}_{k+1}^{i}$ $=$ $\boldsymbol{\nu}_{k}^{i}+\frac{\rho}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left(\boldsymbol{x}_{k+1}^{i}-\boldsymbol{x}_{k+1}^{j}\right),$
$\boldsymbol{\xi}_{k+1}^{i}$ $=$ $\boldsymbol{\xi}_{k}^{i}+\frac{\rho}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left(\boldsymbol{w}_{k+1}^{i}-\boldsymbol{w}_{k+1}^{j}\right).$
We conclude that:
$(\boldsymbol{x}^{i},\boldsymbol{w}^{i})_{k+1}$ $=$ $\underset{(\boldsymbol{x}^{i},\boldsymbol{w}^{i})\in\Sigma^{i}}{\arg\min}\left\{f^{i}(\boldsymbol{x}^{i},\boldsymbol{w}^{i})+\boldsymbol{\mu}^{i\top}(\boldsymbol{z}^{i}-\boldsymbol{w}^{i})\right$.
$\quad+\frac{\rho}{2}\left\|\boldsymbol{z}^{i}-\boldsymbol{w}^{i}\right\|_{\mathbb{R}^{m}}^{2}+(\boldsymbol{\nu}_{k}^{i})^{\top}\boldsymbol{x}^{i}$
$\quad+\frac{\rho}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left\|\boldsymbol{x}^{i}-\frac{\boldsymbol{x}_{k}^{i}+\boldsymbol{x}_{k}^{j}}{2}\right\|^{2}$
$\quad+(\boldsymbol{\xi}_{k}^{i})^{\top}\boldsymbol{w}^{i}$
$\left.+\frac{\rho}{2}\sum_{j\in N^{-}(i)\cup N^{+}(i)}\left\|\boldsymbol{w}^{i}-\frac{\boldsymbol{w}_{k}^{i}+\boldsymbol{w}_{k}^{j}}{2}\right\|^{2}\right\}.$
By dividing all dual variables by $\rho$,
$\tilde{\boldsymbol{\mu}}^{\,i}=\frac{\boldsymbol{\mu}^{i}}{\rho},\quad\tilde{\boldsymbol{\nu}}^{\,i}=\frac{\boldsymbol{\nu}^{i}}{\rho},\quad\tilde{\boldsymbol{\xi}}^{\,i}=\frac{\boldsymbol{\xi}^{i}}{\rho},$
and substituting these rescaled forms into the primal iterations yields the iterates (9)-(9e).